### v1

# Exploratory Study into Relations between Cognitive Distortions and Emotional Appraisals

Abstract
In recent years, there has been growing interest in studying cognitive distortions and emotional appraisals from both computational and psychological perspectives. Despite considerable similarities between emotional reappraisal and cognitive reframing as emotion regulation techniques, these concepts have largely been examined in isolation. This research explores the relationship between cognitive distortions and emotional appraisal dimensions, examining their potential connections and relevance for future interdisciplinary studies. Under this pretext, we conduct an exploratory computational study, aimed at investigating the relationship between cognitive distortion and emotional appraisals. We show that the patterns of statistically significant relationships between cognitive distortions and appraisal dimensions vary across different distortion categories, giving rise to distinct appraisal profiles for individual distortion classes. Additionally, we analyze the impact of cognitive restructuring on appraisal dimensions, exemplifying the emotion regulation aspect of cognitive restructuring.
1 Introduction
Understanding the intricate relationship between cognition, emotion, and behavior has long been a central focus of neuroscience and cognitive science. The advent of artificial intelligence (AI) and recent advances in natural language processing (NLP) have enabled computational researchers to contribute to this field by developing models capable of analyzing individuals’ mental and emotional states from textual data. Within this rapidly evolving domain, the automated extraction of cognitive patterns that shape emotions and behaviors has gained significant traction, bridging the gap between psychological theories and computational innovation.
Emotions are expressed through various modalities, including tone of voice, facial expressions, gestures, and language, particularly in written text. This multifaceted expression of emotions has attracted the interest from NLP and computational researchers in recent years Wang et al. (2022); Plaza-del Arco et al. (2024). While discrete emotional states such as anger, joy, and fear are deemed universal and thus form the basis for automated emotion recognition research, a smaller number of studies have explored dimensional models, representing discrete emotions in continuous spaces Plaza-del Arco et al. (2024). Appraisal theories define emotions as responses that arise from an individual’s evaluation of and event’s significance to their personal goals and well-being, emphasizing that the quality and intensity of emotional responses depend on appraisals, which are the subjective interpretations of the situation Moors et al. (2013).
Negative thoughts are a natural part of human experience; however, they can have a more profound impact on individuals with mental disorders, often becoming entrenched, automatic, and emotionally triggering. Cognitive distortions refer to irrationally exaggerated negative assessments of oneself or situations Beck (1963). Cognitive restructuring, also known as cognitive reframing, is a therapeutic intervention designed to encourage a more positive outlook towards situations by addressing these negative thought patterns Clark (2013). This technique involves replacing negative thoughts with more neutral or hopeful “reframed thoughts”, which provide a softer alternative perspective on the situation.
Although cognitive reappraisal and cognitive restructuring focus on different aspects of cognition— reappraisal aims to change the appraisal of specific events, while restructuring addresses broader thinking patterns—they are both emotion regulation methods that target thoughts to influence emotional states. This suggests the potential for a systematic relationship between emotional appraisal dimensions and cognitive distortions. However, to our knowledge, this relationship has not been thoroughly explored. NLP offers a more accessible and efficient means for conducting such exploratory research compared to traditional psychological studies, which requires recruiting human subjects and eliciting relevant information from them.
Our aim in this work is to bring together emotion appraisals and cognitive distortions, to explore the link between these two different but related psychological constructs. We begin by training appraisal prediction models to perform automated appraisal annotations on a dataset annotated with cognitive distortions enabling a combined analysis of both constructs. We analyze the distribution of appraisal values for each distortion-appraisal pair individually, and find statistically significant relations between cognitive distortions and appraisal dimensions, suggesting that different distortion patterns may exhibit distinct appraisal profiles. Finally, when comparing the appraisal profiles between original and reframed texts, we observed a considerable positive shift in several appraisal dimensions, further demonstrating the link between the two constructs and supporting the need for combined study of the two areas.
2 Appraisal Modeling
Figure 1: Overview of the appraisal prediction model.
In our attempt to analyze the relationship between emotional appraisals and cognitive distortions, we require both appraisal and distortion labels for the same text inputs. No such dataset with both labels is currently available. Therefore, we elected to perform automated data annotation in order to generate the desired labels. Although there are datasets that have been annotated for appraisals Troiano et al. (2023), these are collected from neutral sources (since cognitive appraisal is a normative phenomenon of all emotions, functional and dysfunctional), and thus these texts are not likely to contain too many cognitive distortions. To verify this, we conducted a preliminary experiment and applied a trained cognitive distortion prediction model to the appraisal-annnotated dataset Troiano et al. (2023), which resulted in approximately 80% data points assigned to the “no distortion” class, confirming our assumptions. Consequently, the alternate approach was adopted. In particular, we train appraisal prediction model on the crowd-enVent dataset Troiano et al. (2023), and apply it to the thinking trap dataset Sharma et al. (2023b). The remainder of this section explains our methodology for training the appraisal prediction model.
2.1 Crowd-enVent Dataset
The crowd-enVent is an emotion and appraisal based corpus of event descriptions collected by Troiano et al. (2023) as part of their research on emotional appraisals. During the data collection process, annotators recalled personal events and annotated them based on their recollection of emotions and feelings they experienced at the time of the event. The dataset contains 6600 event descriptions annotated with 21 appraisal dimensions on a 5-point Likert scale. [^1] Appendix C provides definitions of appraisal dimensions considered in this study. The dataset is available in pre-defined splits of training (4320 entries), validation (1080 entries) and test (1200 entries) sets. [^2] Available from https://www.romanklinger.de/data-sets/crowd-enVent2023.zip. Please refer to the original paper by Troiano et al. (2023) for more details on the dataset.
2.2 Model Architecture
We use a multi-regression model to predict the ratings of all appraisal dimensions simultaneously. Specifically, we adopt the multi-regression model by Milintsevich et al. (2023) who used it for predicting the severity of eight depression symptoms. The original model was a hierarchical model implemented with sentence-transformers to encode longer documents. For our sentence-level prediction task, we forgo of the hierarchical definition of the model and directly use the sentence-level embeddings for final predictions. Furthermore, the prediction head now produces 21 regression outputs, one for each emotional appraisal dimension considered. Figure 1 provides an overview of the proposed model.
2.3 Experimental Setup
We use the S-RoBERTa Base model for encoding the input text, which is combined with the corresponding S-RoBERTa tokenizer [^3] https://huggingface.co. The model is trained using AdamW optimizer with a learning rate of $10^{-5}$ and SmoothL1Loss as the loss function. The network architecture applies a dropout of 0.3, along with layer norm regularization in the regression head. The overall code is a modified version of work by Milintsevich et al. (2023). [^4] https://git.unicaen.fr/kirill.milintsevich/hierarchical-depression-symptom-classifier
2.4 Results
Our model performs on par with the results from Troiano et al. (2023). We use the root-mean-squared error (RMSE) as the evaluation metric and report macro-RMSE averaged over all appraisal dimensions on the test set of 1.36 compared to 1.40 reported in the original paper. Furthermore, Figure 2 compares the performance of the two models for each appraisal dimension, with our model out-performing Troiano et al. (2023) for 13 out of 21 appraisal dimensions. Thus, we deem our trained model good enough to be used for automated appraisal annotation in the remainder of the paper.
Figure 2: Appraisal ratings prediction accuracy in RMSE for both our model and the results reported by Troiano et al. (2023).
3 Cognitive Distortion and Emotional Appraisal
With the aim of analyzing the relationship between appraisals and distortions, we apply the appraisal prediction model trained in the previous section to the Thinking Trap dataset Sharma et al. (2023b), obtaining a dataset with 22 labels per input text (one cognitive distortion label and 21 emotional appraisal ratings). The resulting dataset forms the basis for the analysis discussed in the remainder of the paper. This section provides details on the dataset and explains the statistical methods used to analyze the relationships between cognitive distortions and appraisal dimensions.
3.1 Thinking trap Dataset
The Thinking Trap dataset was collected by Sharma et al. (2023b) as part of cognitive reframing research. The dataset contains entries from the existing Thought Records Dataset Burger et al. (2021), along with additional entries collected through an online survey on the Mental Health America website, constituting 300 entries in total. This data was further annotated by mental health practitioners and clinical psychology graduate students, providing annotations for 15 cognitive distortion labels (14 distortions and one “no distortion” class) addressed in the text along with corresponding reframed thoughts [^5] Please refer to appendix C for distortion definitions..
The original annotation process of the Thinking Trap dataset resulted in a multi-label dataset, with each text associated with one or more distortion labels. Because in this research we are interested in each cognitive distortion category separately, we converted the dataset into a multi-class format by repeating the data points once for each associated distortion label. The resulting dataset contained only 19 data points belonging to the “no distortion” class, amounting to only 1.9% of the total data. Because we wanted to contrast appraisal profiles for texts with and without cognitive distortions, we included 77 data points with the “no distortion” class from an additional dataset also collected by the same authors. The final dataset contains 1036 data points with the class distribution provided in Table  1. For more details on the dataset, please refer to the original work by Sharma et al. (2023b).
Distortion Class # entries
All-or-nothing thinking 99
Blaming 34
Catastrophizing 68
Comparing and Despairing 12
Disqualifying the Positive 40
Emotional reasoning 43
Fortune telling 78
Labeling 102
Magnification 15
Mind reading 71
Negative feeling or emotion 151
Overgeneralization 107
Personalization 98
Should statements 22
Not distorted 96
Table 1: Distribution of the cognitive distortion labels of our version of the Thinking Trap dataset.
3.2 Statistical Analysis
To investigate the relationship between cognitive distortions and appraisal, we analyzed the statistical significance between each distortion category and each appraisal dimension. For each distortion-appraisal pair, we formed two groups of texts: a positive group (p), consisting of texts annotated with the cognitive distortion, and a negative group (n), consisting of texts without the distortion. This grouping allowed us to compare appraisal values in the presence and absence of each cognitive distortion.
We performed an independent statistical analysis for each distortion-appraisal pair to isolate the effect of each distortion on the appraisal dimensions. Specifically, we employed the non-parametric Mann-Whitney U Rank Test Mann and Whitney (1947) to assess differences between the positive and negative groups. Under the null hypothesis, we posited that there would be no difference in appraisal values between the two groups (p=n, where p represents the positive group and n represents the negative group). To account for multiple comparisons across 14 cognitive distortion classes and 21 appraisal dimensions, we applied a Bonferroni correction (Abdi et al., 2007), setting a base p-value of 0.05, which was divided by the number of comparisons, which is 307 (the product of 14 and 21).
(a) No-distortion
(b) Exclusive
Figure 3: Discretized significance plots of distortion-appraisal pairs for two definitions of the negative groups. White cell implies a statistically significant relation between the distortion-appraisal pair, the black cell represents no statistically significant difference.
3.3 Negative groups
One major consideration is the definition of the negative group within this analysis. The relationship between cognitive distortions and emotional appraisal dimensions, as inferred from the Mann-Whitney test, is strongly influenced by how the positive and negative distributions are defined. Although the definition of the positive group p is fixed, the negative group can have different meanings. Therefore, we consider the following three different definitions of the negative group n:
No distortion: in this case, the negative distribution only contains entries without any distortion, i.e., those belonging to the “no distortion” class. This group represents the appraisal profile of texts without cognitive distortions and acts as a global baseline against which we can compare individual distortion profiles.
Exclusive: here, the negative distribution contains entries that do not belong to the given distortion class but belong to other cognitive distortion classes (excluding “no distortion”). By utilizing this negative group, we can identify differences in appraisal values between various distortion classes. This approach enables us to analyze the appraisal profile of a specific distortion in relation to other distortions, rather than comparing it to a global baseline.
All others: in this configuration we combine both settings by including in the negative group all entries that do not belong to the specified distortion class.
3.4 Results
Figure 3 plots the discretized significance values for two definitions of the negative groups considered in our study. [^6] Please refer to Appendix A for remaining plots. In these plots, a cell colored white represents a statistically significant difference in appraisal values between the positive and negative groups, while a black cell indicates no statistically significant difference.
First, we observe notable similarities between the two plots. For instance, both plots indicate a lack of statistical significance for emotional reasoning and magnification across all appraisal dimensions considered. Similarly, unpleasantness, goal support, and not consider dimensions exhibit a lack of statistical significance across all distortion classes. However, we also observe certain differences between the two plots. In the “no distortion” setting (Figure 3(a)), appraisal dimensions like chance responsibility and chance control show a significant correlation with more than half of the cognitive distortions. However, in the “exclusive” setting (Figure 3(b)), these dimensions lack significant correlation with any of the distortions. While these plots reveal significant correlations between cognitive distortions and appraisal dimensions, they do not indicate the direction of strength of these correlations, thus motivating the further analysis conducted in the next section.
4 Distortions and Corresponding Appraisal Profiles
The previous section showed systematic relations between cognitive distortions and emotional appraisals. In this section, we delve deeper into these correlations, examining their nature, and studying specific appraisal profiles associated with each distortion class.
Figure 4: Baseline appraisal profile associated with the “no distortion” class.
Figure 5: Appraisal profiles (relative to the “no distortion” baseline) for all distortion classes.
4.1 Methodology
We begin by defining the “baseline” appraisal profile using the “no distortion” negative group. This choice is motivated by the desire to establish a common baseline that represents the appraisal profile of inputs devoid of any cognitive distortion. Specifically, the baseline profile is determined by calculating the median value for each appraisal dimension using the data from the negative group. Similarly, the appraisal profile for each distortion class is defined by calculating the median value for each appraisal dimension using the corresponding positive group data. Finally, to compare the distortion-specific appraisal profiles with the baseline "no distortion" profile, we subtract the baseline profile from each distortion-specific profile. In this manner, we obtain an appraisal profile for each cognitive distortion that is relative to the baseline profile.
4.2 Results
The baseline profile shown in Figure 4 represents the median appraisal values of event descriptions without any cognitive distortions. Firstly, for most appraisal dimensions the median values cluster around the middle of the scale (between scores 2 and 4). Only four dimensions—-familiarity, predictability, unpleasantness, and attention—exhibit median values above 4, while three dimensions—pleasantness, goal support, and social norms—have median values below 2. We also notice a high median value for unpleasantness stemming from the bias in the Thinking Trap dataset, which primarily focused on negative thoughts and situations.
Figure 5 illustrates appraisal profiles associated with individual distortion classes, relative to the baseline profile. The distortion profiles generally exhibit similar patterns, compared to the baseline profile. The two notable exceptions from others are should statements and comparing and despairing, which display deviations in the dimensions of familiarity, other’s responsibility, and other’s control. Regardless of the similarity of the overall pattern, some cognitive distortions show notable peaks in some appraisal dimensions, such as high self responsibility for the should statements or high other’s responsibility and low self responsibility for blaming.
Note that these plots illustrate the relative differences between the appraisal profiles of cognitive distortions and the baseline profile, but they do not indicate which of the distortion-appraisal relations were statistically significant. The following subsection discusses the appraisal profiles, considering the statistical significance analyses presented in Section 3.
4.3 Discussion
While Figure 3 reveals variations in the statistical significance of appraisal dimension correlations across different settings (presence/absence of distortions), Figure 5 demonstrates that the magnitude and direction of appraisal shifts relative to a non-distorted baseline are broadly similar across most distortion classes. This indicates that the presence of cognitive distortions, rather than their specific type, may be the primary driver of altered emotional experiences.
Furthermore, we observe that the appraisal dimensions of suddenness, unpleasantness, goal support, accept consequences, not consider, and effort exhibit a lack of significant correlation with nearly all distortion classes, as illustrated in Figure 3(a). In Figure 5, these dimensions also demonstrate a relatively balanced distribution of distortion profiles around the baseline. In contrast, the appraisal dimensions of chance responsibility, other’s responsibility, other’s control, and chance control show the highest number of significant correlations with distortion classes, as indicated in Figure 3. Additionally, these dimensions exhibit highly polarized values in their distortion profiles, as seen in Figure 5. This correlation between statistical significance and profile polarization suggests that most cognitive distortions exert similar effects on some appraisal dimensions.
Finally, we illustrate specific distortion-appraisal correlations from Figure 3 that align with established psychological principles. To this end, Figure 6 depicts the appraisal profiles for two distortion classes: mind reading and catastrophizing. In the case of mind reading, the observed appraisal values for responsibility (namely, self responsibility, other’s responsibility, and chance responsibility) and control (namely, self control, other’s control, and chance control) are consistent with the distortion’s underlying mechanism. Mind reading, by definition, involves assuming knowledge of another person’s thoughts and intentions, thereby implicitly attributing greater responsibility and control to that other person rather than to oneself or to chance. Conversely, catastrophizing exhibits a negative correlation with the familiarity and accept consequences dimensions. This is psychologically plausible, as an individual’s lack of familiarity with a situation would likely amplify feelings of uncertainty and uncontrollability, making the potential outcomes seem more catastrophic. Furthermore, a reduced ability to accept consequences would logically exacerbate the perceived severity of potential negative outcomes, thus fueling catastrophic thinking.
Figure 6: Appraisal profiles (relative to the baseline) for selected distortion classes.
5 Cognitive Reframing and Emotional Regulation
Figure 7: The shift in appraisal profiles after cognitive reframing. Positive values mean that median appraisal value increase after the reframing, while negative values mean indicate the decrease.
In the final analysis of this research, we examine the impact of cognitive restructuring on appraisal profiles. Cognitive restructuring aims to regulate an individual’s emotional state. Therefore, a significant change in appraisal profiles is expected following the restructuring process. Specifically, we anticipate a positive shift in the appraisal profiles as a result of cognitive restructuring.
5.1 Dataset and Methodology
The Thinking Trap dataset, which formed the basis of our analysis thus far, also provides reframings of the input text. Annotators were asked to write reframes that are rational, specific, readable and actionable. Appraisal labels were generated for the reframes using automated annotation, employing the same prediction model trained in Section 1. Then, based on the reframed inputs, appraisal profiles for different distortion classes were generated using the same methodology detailed in Section 4.1.
5.2 Inferences and Interpretations
Figure 7 plots the difference between appraisal profiles generated from the reframed inputs and the original profiles shown in Figure 5. The plots illustrate a considerable increase in values for pleasantness, while showcasing a decrease in values for dimensions such as unpleasantness and not consider. The observed increase in pleasantness, along with the decrease in unpleasantness, indicates a positive shift in people’s perception of the situations. We also observe a decrease in the appraisal dimension not consider across all distortion classes, reflecting an increased willingness to engage with, rather than avoid, situations. Overall, the changes in appraisal dimensions plotted in Figure  7 provide evidence that cognitive reframing is associated with a potential positive shift in emotional appraisal. This finding further strengthens the proposed link between cognitive distortions and emotional appraisals, and the necessity of studying these concepts together. Finally, this positive shift in appraisal profiles for reframed texts also supports the validity of the automated appraisal prediction model used in our study.
6 Related Work
Although automatic prediction of emotion appraisals has been less studied than predicting discrete emotions, in recent years, several works have emerged in this direction. Several papers have contributed datasets annotated with different sets of appraisal dimensions Hofmann et al. (2020); Troiano et al. (2022, 2023); Zhan et al. (2023). Few experiments have been presented to demonstrate the utility of adopting NLP models to predict the appraisal values based on text, training CNN-based neural classifier Hofmann et al. (2020), fine-tuning RoBERTa-based models Wegge et al. (2022); Troiano et al. (2023), or prompting large language models (LLMs) Zhan et al. (2023).
In relation to cognitive distortions, similarly, few datasets annotated with cognitive distortion categories have been published Shreevastava and Foltz (2021); Wang et al. (2023); Sharma et al. (2023b) with various classification approaches adopted to predict the cognitive distortions from text, using both fine-tuned BERT-based models Tauscher et al. (2023); Maddela et al. (2023) and recently also increasingly prompting LLMs Chen et al. (2023); Lim et al. (2024). Furthermore, NLP researchers have developed methods for a variety of reframing tasks including sentiment and empathy writing Reif et al. (2022); Sharma et al. (2023a), positive reframing Ziems et al. (2022); Goel et al. (2024); Jia et al. (2025), and cognitive restructuring Sharma et al. (2023b); Maddela et al. (2023); Zhan et al. (2024); Xiao et al. (2024).
7 Conclusion
Since both cognitive restructuring and emotional reappraisal serve as emotion regulation strategies, this paper explored the connection between the two constructs from a computational standpoint. As a first step, we automatically annotated appraisal ratings on a dataset of cognitive distortions, producing a new dataset supporting combined analysis. Our analysis at the distortion-appraisal pair level revealed statistically significant relations between cognitive distortions and emotional appraisal dimensions, demonstrating systematic links between the two constructs. By constructing appraisal profiles for individual cognitive distortions and comparing them to a “no distortion” baseline, our analysis showed similar patterns across distortion profiles, indicating a clear distinction between profiles with and without cognitive distortions. Analyzing the impact of cognitive restructuring on the appraisal profiles revealed a shift towards appraisal values indicative of a more positive interpretation of the situations, consistent with the established definition of cognitive reframing. It is our hope that these preliminary results demonstrate the existence of a relationship between cognitive distortions and emotional appraisal dimensions, and illustrate the potential benefits of jointly studying these constructs and motivating further computational research in this area.
Limitations
While we believe this research represents first steps toward understanding correlations between cognitive distortions and emotional appraisals, some concerns remain. A primary issue in computational mental health research is the quality of available data and annotations. The Thinking Trap dataset also suffers from the same problem: manual examination reveals instances with incorrect distortion labels. This classification error stems, in part, from the subjective nature of the task, leading to inconsistencies in labeling even among experienced psychologists. Another concern with this dataset is the length of the input texts. Classifying cognitive distortions or any mental health related aspects based on such short texts is unrealistic and contributes to noise in the data. Some examples illustrating this issue are included in Appendix B.
Another major limitation of this work is that it is exploratory research conducted from a computational standpoint, which lacks the in-depth considerations and reasoning from a psychological perspective. This research needs to be complemented by detailed psychological studies, but this is beyond the scope of this paper as well as our knowledge and expertise.
Using automatic textual analysis to make inferences about people’s mental states has ethical implications. In this work, we use an existing publicly available dataset whose authors reported considering the ethical aspects of their data collection and annotation and also sought approval for their procedures from their institution’s review board (see (Sharma et al., 2023b) Section 4.3). Also, in our study, we are not developing systems for making predictions about individual persons, but rather studying the general patterns over groups.
Appendix
Appendix A Statistical analyses
Since the Thinking Trap dataset ensured a uniform distribution of data points across different distortion classes (Table 1), the appraisal values of texts from cognitive distortion categories outweigh those from the “no distortion” class, resulting in identical plots for the “Exclusive” (Figure 8(c)) and “All others” (Figure 8(a)) settings.
(a) All others
(b) No-distortion
(c) Exclusive
Figure 8: Significance plot between distortions and appraisals for different definitions of negative distribution
Appendix B Data Limitations and Corresponding Examples
The datasets used in this research also exhibit certain quality issues common to most datasets in the field. In the Thinking Trap dataset, some input thoughts are extremely short, making it difficult to assess them due to insufficient information. Some examples of such cases are provided below:
  • I had a breakup, I am the cause of the breakup.
  • I gained weight, I feel like I need to die to be happy.
  • My diet is not working, I feel like a failure.
While these texts may provide some hints about potential cognitive distortions, both models and humans would struggle to accurately assess 21 different emotional appraisal dimensions.
Appendix C Label definitions
C.1 Emotion Appraisals
See Table 2 for the definitions of the 21 emotion appraisal dimensions used in this study.
Dimension Definition
Suddenness The event was sudden or abrupt
Familiarity The event was familiar
Event predictability I could have predicted the occurrence of the event
Pleasantness The event was pleasant
Unpleasantness The event was unpleasant
Goal relevance I expected the event to have important consequences for me
Situational responsibility The event was caused by chance, special circumstances, or natural forces
Self responsibility The event was caused by my own behavior
Others responsibility The event was caused by somebody else’s behavior
Anticipated consequence I anticipated the consequences of the event
Goal support I expected positive consequences for me
Urgency The event required an immediate response
Self control I was able to influence what was going on during the event
Others control Someone other than me was influencing what was going on
Chance control The situation was the result of outside influences of which nobody had control
Consequence acceptance I anticipated that I would easily live with the unavoidable consequences of the event
Internal standards The event clashed with my standards and ideals
External standards The actions that produced the event violated laws or socially accepted norms
Attention I had to pay attention to the situation
Not consider I tried to shut the situation out of my mind
Effort The situation required me a great deal of energy to deal with it
Table 2: List of appraisal dimensions considered in this research.
C.2 Cognitive Distortions
See Table 3 for the definitions of the 14 cognitive distortion categories used in this study.
Dimension Definition
Emotional reasoning Treating your feelings like facts.
Overgeneralization Jumping to conclusions based on one experience.
Disqualifying the positive When something good happens, you ignore it or think it does not count.
Mind reading Assuming that you know what someone else is thinking.
Labeling Defining a person based on one action or characterstic.
Catastrophizing Focusing on the worst/case scenario.
Personalizing Taking things personally, or making them about you
All-or-nothing thinking Thinking in extremes.
Fortune telling Trying to predict the future. Focusing on one possibility and ignoring the other, more likely outcome
Negative feeling and emotion Getting "stuck" on a distressing thought, emotion, or belief.
Blaming Giving away your own power to other people.
Magnification Exaggerating certain aspects of yourself, other people, or a situation while often simultaneously downplaying others.
Comparing and despairing Comparing your worst to someone else’s best.
Should statements Setting unrealistic expectations of yourself.
Table 3: List of cognitive distortions considered in this research.